OpenStudy (anonymous):

what is the sum of these mixed numbers ? 4 1/2 + 5 1/4 =?

jhonyy9 (jhonyy9):

9/2+21/4=(18+21)/4=39/4

jhonyy9 (jhonyy9):

9 3/4

jhonyy9 (jhonyy9):

or you can assum those integers and after those fractions the result will be the same

jhonyy9 (jhonyy9):

4+%=9 1/2 +1/4=3/4

jhonyy9 (jhonyy9):

is it ok ?

jhonyy9 (jhonyy9):

understandably ???

jhonyy9 (jhonyy9):

so if yes please push the good answer button

jhonyy9 (jhonyy9):

you dont like my solutions

jhonyy9 (jhonyy9):

???